Use the following information to answer questions 1 - 13

As of December 31, 2014 assets were $20, liabilities were $12 and paid-in-capital was $1. There was no treasury stock or accumulated other comprehensive income at either the beginning or ending of 2015. During the year revenues were $40, gains (net of losses) were $1, paid-in-capital increased by $2 and the company paid out dividends of $4. At the end of the year, equity was $22 and total liabilities were $20.

1) How much total equity was there on January 1, 2015?

2) How much retained earnings was there at January 1, 2015?

3) How much paid-in-capital was there at December 31, 2015?
Answer #1

(1) Assets = $ 20 and Liabilities = $ 12

Shareholder's Equity = Assets - Liabilities = 20 - 12 = $ 8

Total Equity = Sharholder's Equity = $ 8

(2) Shareholder's Equity = Retained Earnings + Paid-in-Capital + Accumulated Other Comprehensive Income + Treasury Stock

8 = Retained Earnings + 1 + 0 + 0

Retained Earnings = $ 7

(3) During the year the paid-in-capital increased by $ 2 and hence the final paid in capital will be (2+1) = $ 3 at the end of 2015.